OpenStudy (michelle41):

for all non zero r,t and z values, 16r to the 3rd power t the the 5th power divided by -4rt to the 3rd power z squared

OpenStudy (id21):

\[\frac{16r^3t^5}{-4rt^3z^2}\] is it right? Then you can get\[-\frac{4r^2t^2}{z^2}=-(\frac{2rt}{z})^2\]